Cary is 4 years older than Dan. Cary is 4 years older than Find the age of each man now.
Answer by Cromlix(4381)

Cary is 4 years older than Dan
C = 4 + D
In 3 years the sum of their ages will be 68
(D + 3)+ (C + 3) = 68
D + 3 + C + 3 = 68
Replace C with (4 + D) from
top equation.
D + 3 + (4 + D) + 3 = 68
D + 3 + 4 + D + 3 = 68
Collect like terms.
2D + 10 = 68
2D = 68 - 10
2D = 58
D = 29
Currently
Dan is 29 years old.
Cary is 33 years old.
Hope this helps.
:-)